OpenStudy (anonymous):

write the equation of the line in point-intercept form that passes through the points (3,-5) and (9,-2)

OpenStudy (anonymous):

Y+5 = (1/2) (X-3)

OpenStudy (anonymous):

i need the answer in point intercept form, y=mx+b